40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 18 01:14:23 2008, rob, Summary, Computers, EPICS BAD 
    Reply  Thu Sep 18 12:16:01 2008, Yoichi, Update, Computers, EPICS BACK 
       Reply  Thu Sep 18 13:05:05 2008, Yoichi, Update, Computers, EPICS BACK 
Message ID: 964     Entry time: Thu Sep 18 13:05:05 2008     In reply to: 963
Author: Yoichi 
Type: Update 
Category: Computers 
Subject: EPICS BACK 

Quote:

The Y-arm doesn't, because one of the digital filters for the Y-arm lock fails to be loaded to the frontend.
I'm working on it now.


Rob told me that the filter "3^2:20^2" is switched on/off dynamically by the front end code for the LSC.
Therefore, the failure to manually load it was not actually a problem.
The Y-arm did not lock just because the alignment was bad.
Now the Y-arm alignment is ok and the arm locks.
ELOG V3.1.3-